OpenStudy (anonymous):

Will award medal Given three points P(1, –1, 0), Q(0, 1, 2) and R(–1, –1, 1), find the distance from the point Q to the line passing through the points P and R.

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

@Zarkon

OpenStudy (mathmate):

Have you done cross product,magnitudes,... with vectors? @pmkat14

OpenStudy (anonymous):

yes. so what im kinda confused on is do i need to do a projection then a component

OpenStudy (mathmate):

There are different ways of doing it. I propose to find the area of triangle PQR, and divide by the base PR, using vectors.

OpenStudy (anonymous):

so i have PQ<-1,2,2> PR<-2,0,1>. where PR=v. and a. while PQ=b

OpenStudy (anonymous):

then i find the equation of the line as l=(1,-1,0)+t<-2,0,1>

OpenStudy (anonymous):

I saw online that i need to take the unit vector of v. but i dont know where that reasoning comes from

OpenStudy (mathmate):

That's one way. The area way is easier. Twice the area of triangle PQR is the area of the parallelogram given by PQ x RQ. The length of the base is |PR| (magnitude) so the distance d of Q from PR is just |PQ x RQ| / |PR| I'll try to draw it.

OpenStudy (mathmate):



OpenStudy (mathmate):

Examle, |PR|=\(\sqrt{2^2+0^2+(-1)^2}=\sqrt 5\)

OpenStudy (mathmate):

Your line is good.

OpenStudy (mathmate):

The other way is to find vector QP, then subtract vectorially the projection of QP onto PR, which gives the perpendicular of Q onto PR. This is ok too!

OpenStudy (mathmate):